2x/3-5=5x/4+9
solve for x=
2x over 3 as a fraction - 5= 5x over 4 as a fraction +9

Answer by Juhi(44)

2x/3 - 5x/4 = 9 + 5
-7x/12 = 14
-x = -14 * 12/7
= -24